# Universality of high-dimensional scaling limits of stochastic gradient descent

### v1

@@ Abstract.
We consider statistical tasks in high dimensions whose loss depends on the data only through its projection into a fixed-dimensional subspace spanned by the parameter vectors and certain ground truth vectors. This includes classifying mixture distributions with cross-entropy loss with one and two-layer networks, and learning single and multi-index models with one and two-layer networks. When the data is drawn from an isotropic Gaussian mixture distribution, it is known that the evolution of a finite family of summary statistics under stochastic gradient descent converges to an autonomous ordinary differential equation (ODE), as the dimension and sample size go to $\infty$ and the step size goes to 0 commensurately. Our main result is that these ODE limits are universal in that this convergence occurs even when the data is drawn from mixtures of product measures provided the first two moments match the corresponding Gaussian distribution and the initialization and ground truth vectors are sufficiently coordinate-delocalized. We complement this by proving two corresponding non-universality results. We provide a simple example where the ODE limits are non-universal if the initialization is coordinate aligned. We also show that the stochastic differential equation limits arising as fluctuations of the summary statistics around their ODE’s fixed points are not universal.
@@ 1. Introduction
Since the dimension is diverging with the number of samples, a trajectory-wise limit theory of SGD does not naively make sense. However, for projective models with isotropic Gaussian (or isotropic Gaussian mixture) data distributions, rotation invariance means the law of the loss only depends on the point in parameter space through its Gram matrix and its projections into ground truth vectors (and possibly some additional parameters such as second-layer weights for neural networks). Thus, in these Gaussian projective models, one can make sense of high-dimensional limits by focusing on the evolution of a fixed-dimensional set of “summary statistics” like the Gram matrix and inner products with ground truth vectors. Indeed, this reducibility of the law of the loss has been leveraged to great effect in e.g., [41, 42, 45, 46, 47, 26, 9, 28, 4] to develop high-dimensional limit theorems for SGD (and its variants).
@@
However, when the data distribution is not isotropic Gaussian, this reducibility of the dynamics breaks and makes the problem more challenging. The works [37, 17, 14] have delved into how the dynamics evolve under Gaussian noise but with a non-isotropic covariance profile. In non-Gaussian settings, high-dimensional limit theorems for SGD are even rarer. Universal high-dimensional dynamics has been established in the past decade for Langevin dynamics of spin glasses in [22, 21] and approximate message passing schemes in [3, 16]. Essential to those papers was the Lipschitz dependence on the data entries. Similarly, the data enters at most quadratically into linear regression and online independent component analysis (ICA), where the analyses of [2, 30, 48] did not use Gaussianity. By contrast, for general projective models, the loss function depends on low-dimensional projections of the data in an arbitrarily non-linear fashion. The natural question to study is: To what extent do the limits derived in the Gaussian setting hold for SGD trained on non-Gaussian mixture distribution with the same class means, mixture weights, and in-class covariance?
@@
  • (Ballistic phase) In Theorem 1.3, we show that in the ballistic phase, the ODE scaling limit for the summary statistics is universal and only depends on the first two moments of the mixture components, as long as the initialization and class means are coordinate-delocalized. By this we mean that each entry is $O(d^{-1/2+\epsilon})$, as holds for (normalized) i.i.d. vectors (see Definition  1.2). However, this universality breaks when the initialization is localized on a few coordinates, even in well-studied models such as phase retrieval (see Section 1.2.2).
  • (Diffusive phase) In Theorem 1.7, we show that universality generically does not hold for the fluctuations of the algorithm about fixed points of the ballistic dynamics. Our counter example is a simple single-index model with sub-Gaussian data distribution and coordinate-delocalized initialization.
The main idea of the ballistic universality result is as follows. For projective models, the evolution of the summary statistics is governed by expectations of functions of low-dimensional projections of the data, specifically in the directions of the parameter vectors (and possibly certain ground truth vectors). We show that such expectations are quantitatively close to their Gaussian equivalent provided the projections are in coordinate-delocalized directions. A key technical step is then to ensure that if the initialization is coordinate-delocalized, then SGD remains in a delocalized region for all $O(d)$ time scales. We describe the argument in more detail in Section 1.2.1.
@@ 1.2.1. Proof outline
The main technical work is to show that the evolution of summary statistics under any other noise distribution stays within $o(1)$ of the Gaussian equivalent on all linear timescales. To this end, we leverage the structure of projective models as follows. By standard martingale concentration arguments, the evolution of summary statistics under SGD is governed by their drifts. These drifts are given by expectations of functions of the inner products $\bm{\theta}^{\top}X$. If these inner products satisfy a multivariate central limit theorem (CLT) under the noise distribution $\nu$ (together with some quantitative error rates), then we can treat them like the Gaussian case. The Lyapunov CLT suggests that this holds when the parameter vector $\bm{\theta}$ is coordinate-delocalized, and $Y$ has i.i.d. entries. We need a quantitative CLT of this type, which we derive in Section 3.
Our work is then to establish that for SGD with general $\nu$ noise distribution, if the initialization for $\bm{\theta}$ is coordinate-delocalized, then it remains coordinate-delocalized for all linear timescales with high probability. Specifically, we show that $\|\bm{\theta}\|_{\infty}=O(d^{-1/2+\zeta})$ for $\zeta$ small by a Gronwall argument. This requires sharp quantitative control on the drift of $\langle\theta^{a},e_{i}\rangle$. See Section 4 for this argument. The proof of Theorem 1.3 is then concluded in Section 5.
@@ 1.2.2. Importance of coordinate-delocalized initializations
It is natural to wonder if the coordinate-delocalized condition we require on the initialization is truly needed. When the feature data is isotropic Gaussian, any two initializations with the same limit law for the summary statistics admit the same effective dynamics regardless of how coordinate aligned they are. By contrast, under other i.i.d. noise distributions we consider here, different initializations with the same limiting summary statistics can admit different ballistic dynamics in the limit. We demonstrate this in the well-known phase retrieval problem in Lemma 5.2 below, where coordinate-aligned initializations break ballistic universality.
@@ 1.3. Examples
This is clearly a projective model: The loss function (1.8) is expressible as a function $\psi(\bm{\theta}^{\top}X,y)$ for a smooth function $\psi$ with at most linear growth at infinity, and Assumption 1 holds with $q=1$. When the noise distribution is Gaussian, the effective dynamics for the family of summary statistics $\mathbf{G}_{\lfloor t\delta^{-1}\rfloor}=(\bm{\theta}_{\lfloor t\delta^{-1}\rfloor},\bm{\mu})^{\top}(\bm{\theta}_{\lfloor t\delta^{-1}\rfloor},\bm{\mu})$ were derived in [7, Theorem 5.7] as a direct application of the results of [9]. Applying Theorem 1.3, we arrive at the following universality result.
@@
For each $c\in[{\mathcal{C}}]$ the parameter $\Theta^{c}=(\theta^{c},w^{c})$ generates the $c$’th one-vs-all classifier, where $\bm{\theta}^{c}\in\mathbb{R}^{d\times k_{1}}$ is the first layer weights $(\theta^{1,c},\ldots,\theta^{k_{1},c})$, and $w^{c}\in\mathbb{R}^{k_{1}}$ is the second layer weights in the $c$’th one-vs-all classifier being trained. The activation $g$ is applied entry-wise.
It is again easily seen that this model is projective, and in the case where $X$ is Gaussian mixture, its ballistic dynamics limits have been derived. Indeed, this was done in [39, 9] in the special XOR case (see also the related setup of [25]), and more generally in [6] where the dynamics were studied in conjunction with the local loss landscape geometry. By verifying Assumption 1, we establish their universality.
@@ Acknowledgments
R.G. thanks Subhabrata Sen for interesting discussions related to this problem. The research of R.G. is supported in part by NSF CAREER grant 2440509 and NSF DMS grant 2246780. A.J. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada Research Chairs program [RGPIN-2020-04597, DGECR-2020-00199,CRC-2022-00142]. Cette recherche a été enterprise grâce, en partie, au soutien financier du Conseil de Recherches en Sciences Naturelles et en Génie du Canada (CRSNG), et du Programme des chaires de recherche du Canada.
@@ 2.3. Effective dynamics for Gaussian projective models
We take variance $\sigma^{2}=1$ w.l.o.g. Other variances can be captured by modifying $\psi$ and $\Lambda$.
@@
Since the entries of the vector $\delta JVJ^{\top}$ are exactly $\delta$ times $\mathbb{E}[\langle\nabla H,\nabla u\rangle^{2}]$ and $\delta=o(1)$, that implies that the limit of $\delta JVJ^{\top}$ is the zero-vector. ∎
@@
where $\partial_{1,a}\psi$ is a function of $\bm{\theta}^{\top}X,\mathbf{G},w,y$; the law of $\langle\mu^{J},\mu^{b}\rangle$ is a dimension independent function of $y$; Since the law of $\bm{\theta}^{\top}X$ is only a function of $\mathbf{G}$, so is its expectation, so the entire first expectation is only a function of $(\mathbf{G},w)$. The same is true of the second expectation, because the joint law of $\langle Z,\theta^{a}\rangle,\langle Z,\mu^{b}\rangle$ is a dimension independent function of $\mathbf{G}$, and third expectation because $\langle\theta^{a},\mu^{b}\rangle$ is an entry of $\mathbf{G}$. The same argument applies to $u=\langle\theta^{a},\theta^{b}\rangle$ and $u=w$.
@@
Next, Lemma 2.10 implies that effective volatility $\bm{\Sigma}$ is identically zero. Applying Theorem 2.7 yields the claimed result. Finally, $\mathbf{h}$ is locally Lipschitz, and in fact differentiable, in the summary statistics $u$ because $\psi\in C^{3}$ with derivatives of at most polynomial growth at infinity. See Lemma 2.8 for a short justification). ∎
@@
Writing this out in terms of $\psi$, we see that by the same arguments as above, all of the terms are $o(d^{3})$ except
@@ 3.1. The family of functions to which the bounds are applied
We turn to the second part. Without loss of generality take $i=1$. In this case we start by writing
@@ 4. The parameter stays coordinate-delocalized under SGD
Consider SGD for loss satisfying Assumption 1 with respect to $\nu$ noise distribution. Fix any $R>0$ and $\zeta<1/8$. Suppose that $\Theta_{0}\in{\mathscr{D}}_{\zeta}\cap B_{R}(0)$ and $\bm{\mu}\in{\mathscr{D}}_{\zeta}$. Then for any $T>0$,
$\mathbb{P}(\exists\ell\in[0,T\delta^{-1}\wedge\tau_{R}]:\Theta_{\ell}\notin{\mathscr{D}}_{\nicefrac{{1}}{{8}}})=o(1)\,.$

### v2

@@ Abstract.
We consider statistical tasks in high dimensions whose loss depends on the data only through its projection into a fixed-dimensional subspace spanned by the parameter vectors and certain ground truth vectors. This includes classifying mixture distributions with cross-entropy loss with one and two-layer networks, and learning single and multi-index models with one and two-layer networks. When the data is drawn from an isotropic Gaussian mixture distribution, it is known that the evolution of a finite family of summary statistics under stochastic gradient descent converges to an autonomous ordinary differential equation (ODE), as the dimension and sample size go to $\infty$ and the step size goes to 0 commensurately. Our main result is that these ODE limits are universal in that this limit is the same whenever the data is drawn from mixtures of arbitrary product distributions whose first two moments match the corresponding Gaussian distribution, provided the initialization and ground truth vectors are coordinate-delocalized. We complement this by proving two corresponding non-universality results. We provide a simple example where the ODE limits are non-universal if the initialization is coordinate aligned. We also show that the stochastic differential equation limits arising as fluctuations of the summary statistics around their ODE’s fixed points are not universal.
@@ 1. Introduction
Since the dimension is diverging with the number of samples, a trajectory-wise limit theory of SGD does not naively make sense. However, for projective models with isotropic Gaussian (or isotropic Gaussian mixture) data distributions, rotation invariance means the law of the loss only depends on the point in parameter space through its Gram matrix and its projections into ground truth vectors (and possibly some additional parameters such as second-layer weights for neural networks). Thus, in these Gaussian projective models, one can make sense of high-dimensional limits by focusing on the evolution of a fixed-dimensional set of “summary statistics” like the Gram matrix and inner products with ground truth vectors. Indeed, this reducibility of the law of the loss has been leveraged to great effect in e.g., [49, 50, 53, 54, 55, 30, 10, 35, 5, 2] to develop high-dimensional limit theorems for SGD (and its variants).
@@
However, when the data distribution is not isotropic Gaussian, this reducibility of the dynamics breaks and makes the problem more challenging. The works [45, 18, 15, 19] have delved into how the dynamics evolve under Gaussian noise but with a non-isotropic covariance profile. In non-Gaussian settings, high-dimensional limit theorems for SGD are even rarer. Universal high-dimensional dynamics has been shown in the past decade for Langevin dynamics of spin glasses in [25, 24] and approximate message passing schemes in [4, 17, 27]. Essential to those papers was the Lipschitz dependence on the data entries. Similarly, the data enters at most quadratically into linear regression and online independent component analysis (ICA), where the analyses of [3, 37, 56] did not use Gaussianity. By contrast, for general projective models, the loss function depends on low-dimensional projections of the data in an arbitrarily non-linear fashion.
In light of the above, the natural question to study is: To what extent do the limits derived in the Gaussian setting hold for the trajectory of summary statistics of online SGD trained on non-Gaussian mixture distribution with the same class means, mixture weights, and in-class covariance? Universality of this type was proposed in [32, 31] where it was argued that this should follow from a certain “Gaussian equivalence property” for the outputs of the first layer weights along training, see Conjecture 1 of [31] (see also Section 1.2.3). We prove both of these results for projective models.
@@
  • (Ballistic phase) In Theorem 1.3, we show that in the ballistic phase, the ODE scaling limit for the summary statistics is universal and only depends on the first two moments of the mixture components, as long as the initialization and class means are coordinate-delocalized. By coordinate-delocalized we mean that each entry is $O(d^{-\frac{1}{2}+\epsilon})$, as holds for (normalized) i.i.d. vectors (see Definition  1.2). However, this universality breaks when the initialization is localized on a few coordinates, even in well-studied models such as phase retrieval (see Section 1.2.2).
  • (Diffusive phase) In Theorem 1.8, we show that universality generically does not hold for the fluctuations of the summary statistics about fixed points of the ballistic dynamics. Our counter example is a simple single-index model with sub-Gaussian data distribution and coordinate-delocalized initialization.
The main idea of the ballistic universality result is as follows. For projective models, the evolution of the summary statistics is governed by expectations of functions of low-dimensional projections of the data, specifically in the directions of the parameter vectors (and possibly certain ground truth vectors). Random low-dimensional projections of high-dimensional product measures are well-known to satisfy central limit theorems. In particular, we show these expectations are quantitatively close to their Gaussian equivalent, uniformly over projections into coordinate-delocalized directions. The key step is then to ensure that if the initialization and ground truth vectors are coordinate-delocalized, then SGD remains in a delocalized region for all $O(d)$ time scales. We describe the argument in more detail in Section 1.2.1.
@@ 1.2.1. Proof outline
The aim is then to show that the evolution of summary statistics under any other noise distribution stays within $o(1)$ of the Gaussian equivalent on all linear timescales. To this end, we leverage the structure of projective models as follows. By standard martingale concentration arguments, the evolution of summary statistics under SGD is governed by their drifts. These drifts are given by expectations of functions of the inner products $\bm{\theta}^{\top}X$. If these inner products satisfy a multivariate central limit theorem (CLT) under the noise distribution $\nu$ (together with some quantitative error rates), then we can treat them like the Gaussian case. The Lyapunov CLT suggests that this holds when the parameter vector $\bm{\theta}$ is coordinate-delocalized, and $Y$ has i.i.d. entries. We need a quantitative CLT of this type, which we derive in Section 3.
Our work to show this is primarly then to establish that for SGD with general $\nu$ noise distribution, if the initialization for $\bm{\theta}$ is coordinate-delocalized, then it remains coordinate-delocalized for all linear timescales with high probability. Specifically, we show that $\|\bm{\theta}\|_{\infty}=O(d^{-1/2+\zeta})$ for $\zeta$ small by a Gronwall argument. This requires sharp quantitative control on the drift of $\langle\theta^{a},e_{i}\rangle$. See Section 4 for this argument. The proof of Theorem 1.3 is then concluded in Section 5.
@@ 1.2.2. Importance of coordinate-delocalized initializations
It is natural to wonder if the coordinate-delocalized condition we require on the initialization is truly needed. When the feature data is isotropic Gaussian, any two initializations with the same limit law for the summary statistics admit the same effective dynamics regardless of how coordinate aligned they are. By contrast, under other i.i.d. noise distributions we consider here, different initializations with the same limiting summary statistics can admit different ballistic dynamics in the limit. We demonstrate this in the well-known phase retrieval problem in Proposition 5.1 below, where coordinate-aligned initializations break ballistic universality.
1.2.3. Relation to the “Deep Gaussian Equivalence Conjecture”
The authors of [32, 31] proposed the universality of the law of the output of the first-layer in teacher-student networks (see Section 1.3). This proposal has received much attention in the learning theory literature. Specifically, Conjecture 1 of [31]—which they called the Deep Gaussian Equivalence Conjecture (DGEC)—had two forms: (a) An active learning form, where it was predicted that after $td$ steps of online SGD for $t>0$, the inner product of the features with the first layer weights, i.e., $\bm{\theta}_{\lfloor td\rfloor}^{\top}X$, are asymptotically Gaussian; and (b) An empirical risk minimizer (ERM) form where the same quantity evaluated at the empirical risk minimizer is predicted to be asymptotically Gaussian. There has been significant progress on (b) [31, 34, 23, 41], from which one can deduce universality of the risk evaluated at the ERM. On our way to the proof of Theorem 1.3, we establish the following precise version of the active learning form (a) of the DGEC. Let $\tau_{R}$ be the exit time of $\Theta_{\ell}$ from the $\ell_{2}$ ball of radius $R$ around 0. For any two probability measures, $\mu,\nu,$ let $d_{C^{3}_{b}}(\mu,\nu)=\sup_{\|f\|_{C^{3}_{b}}\leq 1}|\mathbb{E}_{\mu}f-\mathbb{E}_{\nu}f|$.
Theorem 1.4.
Fix any $\zeta<1/8$ and $c_{\textsc{lr}}>0$. There exists $\varepsilon(\zeta,c_{\textsc{lr}})>0$ such that the following holds for all $\bm{\mu}\in\mathcal{D}_{\zeta}$, and step-size $\delta=\frac{c_{\textsc{lr}}}{d}$. Let $\Theta_{\ell}=(\bm{\theta}_{\ell},w_{\ell})$ be SGD with $\nu$-MM data satisfying Assumption 1, and $\bar{\Theta}_{\ell}=(\bar{\bm{\theta}}_{\ell},\bar{w}_{\ell})$ be SGD with GMM data, both initialized from $\Theta_{0}=\bar{\Theta}_{0}\in\mathcal{D}_{\zeta}$. For any $R>0$, with probability 1o(1)1-o(1) over the realization of SGD, as $d\to\infty$,
$\max_{\ell\leq(\varepsilon d\log d)\wedge\tau_{R}\wedge\bar{\tau}_{R}}d_{C^{3}_{b}}(\mathrm{Law}(\bm{\theta}_{\ell}^{\top}X),\mathrm{Law}(\bm{\theta}_{\ell}^{\top}\bar{X}))\to 0$
where $X,\bar{X}$ are independent data points from $\nu$-MM and GMM respectively.
Proving this DGEC for all linear timescales of online SGD is the bulk of the work for establishing universality of the ballistic limits. However, we emphasize two important caveats: Firstly, if the initialization or ground truth vectors are not sufficiently coordinate-delocalized, the DGEC is not true: see Proposition 5.1. Secondly, the restriction to timescales less than $\tau_{R}\wedge\bar{\tau}_{R}$ is necessary: Due to the level of generality we consider, the norm of the parameters may blow-up in $\Omega(d)$ time. It is easy to see that stronger assumptions on the growth of $\psi$ at infinity, together with positive regularizer $\Lambda$, are enough to guarantee $\tau_{R}\wedge\bar{\tau}_{R}=\omega(d)$ with high probability.
@@ 1.3. Examples
This is clearly a projective model: The loss function (1.8) is expressible as a function $\psi(\bm{\theta}^{\top}X,y)$ for a smooth function $\psi$ with at most linear growth at infinity, and Assumption 1 holds with $q=1$. When the noise distribution is Gaussian, the effective dynamics for the family of summary statistics $\mathbf{G}_{\lfloor t\delta^{-1}\rfloor}=(\bm{\theta}_{\lfloor t\delta^{-1}\rfloor},\bm{\mu})^{\top}(\bm{\theta}_{\lfloor t\delta^{-1}\rfloor},\bm{\mu})$ were derived in [8, Theorem 5.7] as a direct application of the results of [10]. Applying Theorem 1.3, we arrive at the following universality result.
@@
For each $c\in[{\mathcal{C}}]$ the parameter $\Theta^{c}=(\theta^{c},w^{c})$ generates the $c$’th one-vs-all classifier, where $\bm{\theta}^{c}\in\mathbb{R}^{d\times k_{1}}$ is the first layer weights $(\theta^{1,c},\ldots,\theta^{k_{1},c})$, and $w^{c}\in\mathbb{R}^{k_{1}}$ is the second layer weights in the $c$’th one-vs-all classifier being trained. The activation $g$ is applied entry-wise. This is sometimes called a teacher-student network.
It is easy to that this model is projective, and when $X$ is Gaussian mixture, its ballistic dynamics limits have been derived. Indeed, this was done in [47, 10] in the special XOR case (see also the related setup of [29]), and more generally in [7] where the dynamics were studied in conjunction with the local loss landscape geometry. By verifying Assumption 1, we establish their universality.
@@ Acknowledgments
The authors thank Sebastian Goldt and Bruno Loureiro for pointing them to Conjecture 1 of [31] and related literature. R.G. thanks Subhabrata Sen for interesting discussions related to this problem. The research of R.G. is supported in part by NSF CAREER grant 2440509 and NSF DMS grant 2246780. A.J. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada Research Chairs program [RGPIN-2020-04597, DGECR-2020-00199,CRC-2022-00142]. Cette recherche a été enterprise grâce, en partie, au soutien financier du Conseil de Recherches en Sciences Naturelles et en Génie du Canada (CRSNG), et du Programme des chaires de recherche du Canada.
@@ 2.3. Effective dynamics for Gaussian projective models
We take variance $\sigma^{2}=1$ without loss of generality. Other variances can be captured by modifying $\psi$ and $\Lambda$.
@@
The result then follows as $\lVert\delta JVJ^{\top}\rVert=O(\delta\max_{u}\mathbb{E}[\langle\nabla H,\nabla u\rangle^{2}])$ and $\delta=o(1)$.∎
@@
where $\partial_{1,a}\psi$ is a function of $\bm{\theta}^{\top}X,\mathbf{G},w,y$; the law of $\langle\mu^{J},\mu^{b}\rangle$ is a dimension independent function of $y$. Since the law of $\bm{\theta}^{\top}X$ is only a function of $\mathbf{G}$, so is its expectation, so the entire first expectation is only a function of $(\mathbf{G},w)$. The same is true of the second expectation, because the joint law of $\langle Z,\theta^{a}\rangle,\langle Z,\mu^{b}\rangle$ is a dimension independent function of $\mathbf{G}$, and third expectation because $\langle\theta^{a},\mu^{b}\rangle$ is an entry of $\mathbf{G}$. The same argument applies to $u=\langle\theta^{a},\theta^{b}\rangle$ and $u=w$.
@@
Next, Lemma 2.10 implies that effective volatility $\bm{\Sigma}$ is identically zero. Applying Theorem 2.7 yields the claimed result. Finally, $\mathbf{h}$ is locally Lipschitz, and in fact differentiable, in the summary statistics $u$ because $\psi\in C^{3}$ with derivatives in $\mathcal{F}_{q}$. See Lemma 2.8 for a short justification. ∎
@@
Writing this out in terms of $\psi$, we see that, by the same arguments as above, all of the terms are $o(d^{3})$ except
@@ 3.1. The family of functions to which the bounds are applied
We turn to the second part. Without loss of generality take $i=1$. Start by writing
@@ 4. The parameter stays coordinate-delocalized under SGD
Consider SGD for loss satisfying Assumption 1 with respect to $\nu$ noise distribution. For $\zeta<1/8$, there exists $\varepsilon>0$ such that for any $R>0$ the following holds: If $\Theta_{0}\in{\mathscr{D}}_{\zeta}\cap B_{R}(0)$ and $\bm{\mu}\in{\mathscr{D}}_{\zeta}$, then
$\mathbb{P}(\exists\ell\in[0,(\varepsilon\delta^{-1}\log d)\wedge\tau_{R}]:\Theta_{\ell}\notin{\mathscr{D}}_{\nicefrac{{1}}{{8}}})=o(1)\,.$
Also in this section, we deduce Theorem 1.4 as a consequence of Theorem 4.1.